Write conversion factors (or coefficient ratio) for the following reaction: 4NH3 + 5O2 → 4NO +...

  1. Write conversion factors (or coefficient ratio) for the following reaction:

4NH3 + 5O2 → 4NO + 6H2O

  1. The conversion factor that converts mole of NH3 to mole of H­2O

       mole of NH3 x ____________  =  mole of H­2O

  1. The conversion factor that converts mole of NH3 to mole of NO

mole of NH3 x ____________  =  mole of NO

  1. The conversion factor that converts mole of NH3 to mole of O­2

mole of NH3 x ____________=  mole of O­2

  1. The conversion factor that converts mole of O­2 to mole of  NO

             mole of O2 x ___________=  mole of  NO
Answer #1

(a) 4 moles of NH3 produces 6 moles of H2O so conversion factor = 6/4 = 1.5

(b) 4 moles of NH3 produces 4 moles of NO, hence conversion factor is = 4/4 = 1

(c) 4 moles of NH3 require 5 moles of O2 so conversion factor of NH3 to O2 = 5/4 = 1.25

(d) 5 moles O2 produces 4 moles of NO, so conversion factor = 4/5 = 0.8